BRICS Notes Series, 2000

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-00-8] Abstract, PostScript, PDF.

Anders Mller and Michael I. Schwartzbach. The XML Revolution. December 2000. 149 pp.

[NS-00-7] Abstract, PostScript, PDF, DVI.

Nils Klarlund, Anders Mller, and Michael I. Schwartzbach. Document Structure Description 1.0. December 2000. 40 pp.

[NS-00-6] Abstract, PostScript, PDF.

Peter D. Mosses and Hermano Perrelli de Moura, editors. Proceedings of the Third International Workshop on Action Seman-

tics, AS 2000, (Recife, Brazil, May 15–16, 2000), August 2000. viii+148 pp.

[NS-00-5] Abstract, PostScript, PDF.

Claus Brabrand. <bigwig> Version 1.3 — Tutorial. September 2000. ii+92 pp.

[NS-00-4] Abstract, PostScript, PDF.

Claus Brabrand. <bigwig> Version 1.3 — Reference Manual. September 2000. ii+56 pp.

[NS-00-3] Abstract, PostScript, PDF.

Patrick Cousot, Eric Goubault, Jeremy Gunawardena, Maurice Her-

lihy, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topol-

ogy in Concurrency Theory, GETCO ’00, (State College, USA, August 21, 2000), August 2000.

vi+116 pp.

[NS-00-2] Abstract, PostScript, PDF. Luca Aceto and Bjrn Victor, editors. Preliminary Proceedings of the 7th International Workshop on Ex-

pressiveness in Concurrency, EXPRESS ’00, (State College, Pennsylvania, USA, August 21, 2000), August 2000.

vi+130 pp.

[NS-00-1] Abstract, PostScript, PDF. Bernd Grtner. Randomization and Abstraction — Useful Tools for Optimization. February 2000. 106 pp.